\begin{thebibliography}{99}

\bibitem{BGSV2012B}
C. Bachoc, D. C. Gijswijt, A. Schrijver, and F. Vallentin,
Invariant semidefinite programs,
in: Handbook on semidefinite, conic and polynomial optimization (M. F. Anjos and J. B. Lasserre, eds.), Springer, New York, 2012, pp.~219--269;
\arxiv{1007.2905}.

\bibitem{BCT1999}
G. Brassard, R. Cleve, and A. Tapp,
Cost of exactly simulating quantum entanglement with classical communication,
Phys. Rev. Lett. 83 (1999) 1874--1877;
\arxiv{quant-ph/9901035}.

\bibitem{BCW1998}
H. Buhrman, R. Cleve, and A. Widgerson, 
Quantum vs. classical communication and computation,
in: Proceedings of the 30th Annual ACM Symposium on the Theory of Computing, Dallas, TX, USA, 1998, pp.~63--68;
\arxiv{quant-ph/9802040}.

\bibitem{Cameron2010EJC}
P. J. Cameron,
Problems from CGCS Luminy, May 2007,
European J. Combin. 31 (2010) 644--648.

\bibitem{Delsarte1973}
P. Delsarte,
An algebraic approach to the association schemes of coding theory,
Philips Res. Rep. Suppl., No.~10, 1973.

\bibitem{Fine1947AMM}
N. J. Fine,
Binomial coefficients modulo a prime,
Amer. Math. Monthly 54 (1947) 589--592.

\bibitem{Frankl1986C}
P. Frankl,
Orthogonal vectors in the $n$-dimensional cube and codes with missing distances,
Combinatorica 6 (1986) 279--285.

\bibitem{FR1987}
P. Frankl and V. R\"{o}dl,
Forbidden intersections,
Trans. Amer. Math. Soc. 300 (1987) 259--286.

\bibitem{Galliard2001D}
V. Galliard,
Classical pseudo-telepathy and colouring graphs,
diploma thesis, ETH Zurich, 2001;
available at \url{http://math.galliard.ch/Cryptography/Papers/PseudoTelepathy/SimulationOfEntanglement.pdf} .

\bibitem{GTW2003P}
V. Galliard, A. Tapp, and S. Wolf,
The impossibility of pseudo-telepathy without quantum entanglement,
in: Proceedings 2003 IEEE International Symposium on Information Theory, Yokohama, Japan, 2003;
\arxiv{quant-ph/0211011}.

\bibitem{GMS2012IEEE}
D. C. Gijswijt, H. D. Mittelmann, and A. Schrijver,
Semidefinite code bounds based on quadruple distances,
IEEE Trans. Inform. Theory 58 (2012) 2697--2705;
\arxiv{1005.4959}.

\bibitem{GN2008SIAM}
C. D. Godsil and M. W. Newman,
Coloring an orthogonality graph,
SIAM J. Discrete Math. 22 (2008) 683--692;
\arxiv{math/0509151}.

\bibitem{KP2007EJC}
E. de Klerk and D. V. Pasechnik,
A note on the stability number of an orthogonality graph,
European J. Combin. 28 (2007) 1971--1979;
\arxiv{math/0505038}.

\bibitem{Laurent2007MP}
M. Laurent,
Strengthened semidefinite programming bounds for codes,
Math. Program. 109 (2007) 239--261.

\bibitem{Newman2004D}
M. W. Newman,
Independent sets and eigenspaces,
thesis, University of Waterloo, 2004.

\bibitem{Schrijver2005IEEE}
A. Schrijver,
New code upper bounds from the Terwilliger algebra and semidefinite programming,
IEEE Trans. Inform. Theory 51 (2005) 2859--2866.

\end{thebibliography}

\documentclass{amsart}

\usepackage{amssymb,amsthm,mathrsfs}
\usepackage{hyperref}

\newcommand{\arxiv}[1]{arXiv:\linebreak[0]\href{http://arxiv.org/abs/#1}{#1}}

\newtheorem*{thm}{Theorem}
\theoremstyle{definition}
\newtheorem*{prob}{Problem}

\begin{document}

\title[The independence number of the orthogonality graph]{The Independence Number of the Orthogonality Graph in Dimension $2^k$}

\author{Ferdinand Ihringer}
\address{Department of Mathematics: Analysis, Logic and Discrete Mathematics, Ghent University, Belgium}
\thanks{The first author is supported by a postdoctoral fellowship of the Research Foundation --- Flanders (FWO)}
\email{ferdinand.ihringer@ugent.be}

\author{Hajime Tanaka}
\address{\href{http://www.math.is.tohoku.ac.jp/}{Research Center for Pure and Applied Mathematics}, Graduate School of Information Sciences, Tohoku University, Japan}
\thanks{The second author is supported by JSPS KAKENHI Grant Number JP17K05156.}
\email{htanaka@tohoku.ac.jp}

\begin{abstract}
  We determine the independence number of the orthogonality graph on $2^k$-dimensional hypercubes.
  This answers a question by Galliard from 2001 which is motivated by a problem in quantum information theory.
  Our method is a modification of a rank argument due to Frankl 
  who showed the analogous result for $4p^k$-dimensional hypercubes, where $p$ is an odd prime.
\end{abstract}

\maketitle

\section{Introduction}

The \textit{orthogonality graph} $\Omega_n$ has the elements of $\{ -1, 1 \}^n$ as vertices, and two vertices
are adjacent if they are orthogonal, in other words, if their Hamming distance is $n/2$.
The graph $\Omega_n$ occurs naturally when comparing classical and quantum communication \cite{BCW1998}.
In particular, for $n=2^k$ the cost of simulating a specific quantum entanglement on $k$ qubits 
can be reduced to determining the chromatic number $\chi(\Omega_{n})$ of $\Omega_n$ \cite{BCT1999,Galliard2001D}.
The graph $\Omega_n$ is edgeless if $n$ is odd, and is bipartite if $n\equiv 2 \pmod{4}$.
For $n\equiv 0 \pmod{4}$, Frankl \cite{Frankl1986C} and Galliard \cite{Galliard2001D} constructed an independent set of $\Omega_n$ of size
\begin{equation*}\label{pin}
	a_n := 4 \sum_{i=0}^{n/4-1} \binom{n-1}{i},
\end{equation*}
and Galliard \cite{Galliard2001D} asked in 2001 if this is the independence number $\alpha(\Omega_n)$ of $\Omega_n$ when $n=2^k$, $k\geqslant 2$.
Newman \cite{Newman2004D} and, according to \cite[p.~275, Remark]{FR1987}, Frankl conjectured that this holds whenever $n\equiv 0\pmod{4}$.
See also \cite{Cameron2010EJC}. 
Frankl \cite{Frankl1986C} already showed the conjecture in 1986 for all $n=4p^k$ for $k \geqslant 1$, where $p$ is an odd prime.
De Klerk and Pasechnik \cite{KP2007EJC} proved the conjecture for $n=16$, i.e., that $\alpha(\Omega_{16})=2304$, 
using Schrijver's semidefinite programming bound \cite{Schrijver2005IEEE}. Furthermore,
Frankl and R\"{o}dl \cite{FR1987} showed that $\alpha(\Omega_{n}) < 1.99^n$ if $n\equiv 0\pmod{4}$.
In this note, we apply Frankl's method from \cite{Frankl1986C} to show the following:
\begin{thm}
Let $n = 2^k$ for some $k \geqslant 2$.
Then $\alpha(\Omega_n)=a_n$.
\end{thm}
Together with the discussion in \cite[Section 5.5]{Galliard2001D}, that is using $\chi(\Omega_n) \geqslant 2^n/\alpha(\Omega_n)$, 
our result implies an explicit version of Theorem 4 in \cite{BCT1999}.
Finding such an explicit result is one motivation for Galliard's work.
See also \cite{GTW2003P,GN2008SIAM}.

\section{Proof of the Theorem}

Let $A_j$ be the $0$-$1$-matrix indexed by the vertices of the hypercube $Q_n = \{ -1, 1 \}^n$ with $(A_j)_{xy} = 1$ if 
$x$ and $y$ have Hamming distance $j$.
The matrices $A_j$ have $n+1$ common  eigenspaces $V_0, V_1, \dots, V_n$, and
in the usual ordering of the eigenspaces the eigenvalue of $A_j$ with respect to $V_i$
is given by the Krawtchouk polynomial (see \cite[Theorem 4.2]{Delsarte1973})
\begin{align*}
 	K_j(i)=K_j(i; n) := \sum_{h=0}^j (-1)^h \binom{i}{h} \binom{n-i}{j-h}.
\end{align*}
It is known that the orthogonal projection matrix $E_i$ onto $V_i$ has
the entry $(E_i)_{xy} = 2^{-n} K_i(j)$ if $x$ and $y$ are at Hamming distance $j$ \cite[Theorem 4.2]{Delsarte1973}, so that we have in particular $\operatorname{rank} E_i=\operatorname{trace} E_i=K_i(0)= \binom{n}{i}$.
The $(n+1)$-dimensional matrix algebra
spanned by $A_0=I, A_1, \dots, A_n$ is called the \textit{Bose--Mesner algebra} of $Q_n$.

Assume now that $n=2^k$, $k\geqslant 3$.
(The result is trivial if $k=2$.)
Let $C$ be an independent set of $\Omega_{2^k}$, and let $C_{\mathrm{even}}^{\pm},C_{\mathrm{odd}}^{\pm} \subseteq \{ -1, 1 \}^{2^k-1}$ be as in \cite{Frankl1986C}:
$C_{\mathrm{even}}^+$ is given by taking all the even-weight elements of $C$ that end with $+1$, followed by truncating at the last coordinate, and the other three are analogous.
Let $C'$ be one of these four families.
Then the Hamming distances in $C'$ are even and unequal to $2^{k-1}$, so they lie in the following set:
\begin{equation}\label{weights}
	\{2s : s=0,1,\dots,2^{k-1}-1,\, s\ne 2^{k-2}\}.
\end{equation}

Below we work with the Bose--Mesner algebra $\mathscr{A}$ of $Q_{2^k-1}$.
For every $M\in\mathscr{A}$, let $\overline{M}$ denote the principal submatrix corresponding to $C'$.
Consider the polynomial
\begin{equation*}
	\varphi(\xi)=\binom{\xi/2-1}{2^{k-2}-1} \in \mathbb{R}[\xi],
\end{equation*}
and expand it in terms of the Krawtchouk polynomials $K_i(\xi)=K_i(\xi;2^k-1)$:
\begin{equation}\label{Krawtchouk expansion}
	\varphi(\xi)=\sum_{i=0}^{2^{k-2}-1} c_i K_i(\xi).
\end{equation}
Let
\begin{equation*}
	X=\sum_{j=0}^{2^k-1}\varphi(j)A_j \in\mathscr{A}.
\end{equation*}
On the one hand, observe that $\overline{X}$ has only integral entries in view of \eqref{weights}, and an easy application of Lucas' theorem (cf.~\cite{Fine1947AMM}) shows moreover that $\overline{X}\equiv \overline{I} \pmod{2}$.
In particular, $\overline{X}$ is invertible.
On the other hand, from \eqref{Krawtchouk expansion} we have
\begin{equation*}
	X=2^{2^k-1}\sum_{i=0}^{2^{k-2}-1} c_i E_i.
\end{equation*}
It follows that
\begin{align*}
	|C'| &= \operatorname{rank} \overline{X} \leqslant \operatorname{rank} X 
	 \leqslant \sum_{i=0}^{2^{k-2}-1} \operatorname{rank} E_i = \sum_{i=0}^{2^{k-2}-1} \binom{2^k-1}{i}.
\end{align*}
As $|C| = |C_{\mathrm{even}}^+| + |C_{\mathrm{even}}^-| + |C_{\mathrm{odd}}^+| + |C_{\mathrm{odd}}^-|$, the theorem follows.

\section{Future Work}

Schrijver's semidefinite programming bound has been extended to hierarchies of upper bounds; see, e.g., \cite{BGSV2012B,Laurent2007MP}.
In view of \cite{KP2007EJC}, it is interesting to investigate if these bounds in turn prove the conjecture for other values of $n$.
One of the referees pointed out to us that using next level in the hierarchy, see \cite{GMS2012IEEE}, yields the correct bound
of $a_{24}=178208$ for the case $n=24$.

\begin{prob}
Prove the conjecture for $n=40$, which is the first open case.
\end{prob}

\paragraph*{Acknowledgements} We thank the anonymous referee for solving the case $n=24$.

\end{document}